# Language Model Council: Benchmarking Foundation Models on Highly Subjective Tasks by Consensus

### v1

@@ Abstract
The rapid advancement of Large Language Models (LLMs) necessitates robust and challenging benchmarks. Leaderboards like Chatbot Arena rank LLMs based on how well their responses align with human preferences. However, many tasks such as those related to emotional intelligence, creative writing, or persuasiveness, are highly subjective and often lack majoritarian human agreement. Judges may have irreconcilable disagreements about what constitutes a better response. To address the challenge of ranking LLMs on highly subjective tasks, we propose a novel benchmarking framework, the Language Model Council (LMC). The LMC operates through a democratic process to: 1) formulate a test set through equal participation, 2) administer the test among council members, and 3) evaluate responses as a collective jury. We deploy a council of 20 newest LLMs on an open-ended emotional intelligence task: responding to interpersonal dilemmas. Our results show that the LMC produces rankings that are more separable, robust, and less biased than those from any individual LLM judge, and is more consistent with a human-established leaderboard compared to other benchmarks.
@@ 1 Introduction
We are experiencing a Cambrian Explosion of Large Language Models (LLMs) that exhibit remarkable and wide-ranging abilities. As LLMs have advanced, evaluating their quality has become increasingly challenging. Publishers of new models often claim superiority based on various benchmarks, citing their rankings on leaderboards like [6, 10, 15, 25, 28].
Conventional LLM evaluations use closed-ended questions (e.g., multi-choice questions like MMLU [20] or TruthfulQA [26]) that can be automatically verified. However, the static nature of benchmarks risks contamination, where models may have been inadvertently exposed to elements of the test datasets during training, thereby skewing evaluation results [31].
@@
Figure 1: Overview of benchmarking by Language Model Council. Using the same LLMs for test set formulation, task completion, and judging offers an equitable way to achieve a ranking by consensus on highly subjective tasks.
Arena-based approaches such as [44, 7, 10, 43, 15, 24] aim to address this limitation by evaluating LLMs on open-ended questions. These methods typically involve two models competing in various tasks to evaluate their abilities. Examples of these tasks include responding to open-ended questions, completing specific tasks, playing skill-based games, or having multi-turn conversations. The outcomes of these battles serve as indicators of each model’s competency and can be determined objectively (NegotiationArena [7]), by a strong judge model like GPT-4 (AlpacaEval [15], Chatbot Arena Hard [24]), by a committee of LLM judges (PoLL [36], Auto Arena [43]), or by real humans (MT-Bench and Chatbot Arena [24]).
But how consistent are human ratings to begin with? Substantial disagreements are common in labeling online comment toxicity, news misinformation, and medical diagnosis, with up to one-third of expert annotators disagreeing on an average example, even after accounting for label noise [18]. In a recent study on persuasiveness, humans expressed different experiences after reading the same argument 89% of the time [16]. On MT-Bench, human ratings agree only 38% of the time on a per-item basis, with 0% unanimity among expert raters [44].
There is growing interest in establishing benchmarks for emotional intelligence (EI) in Large Language Models (LLMs) [28, 37, 32]. However, assessing EI is highly subjective, and the relationship between EI and IQ remains contentious, with studies showing positive, negative, or no correlation [27, 17]. On EmoBench, a handcrafted test for emotional intelligence, only 32.5% of questions receive unanimous agreement, despite its simple, multiple-choice format [32].
To tackle the challenge of evaluating LLMs on highly subjective tasks, we introduce the Language Model Council (LMC). The LMC consists of three stages. First, a test set is formulated with equal participation from all council members. Second, the test set is administered to all council members to complete. Lastly, the responses are evaluated by the council as a collective jury.
Our contributions are as follows:
@@
    We propose the LMC as a flexible evaluation framework for ranking LLM agents on subjective tasks in a democratic fashion. We rank 20 state-of-the-art LLMs based on an open-ended emotional intelligence task that is more separable, robust, and less biased than using any individual LLM judge. Qwen-1.5-110B emerges as the leader, surpassing GPT-4o in second place.
@@
    We define and quantify key measures of LLM judging quality in an ensemble setting such as consistency, agreement, separability, affinity, and bias. We identify claude-3-opus and mistral-large as the most effective judges.
@@
    We run a human study to determine a parallel leaderboard on this EI task for a subset of 9 LLMs. We find that the LMC’s ranking is significantly more correlated with the human-established rankings compared to other benchmarks.
  • 4.
    We use Monte-Carlo simulations to measure ranking stability and robustness to adversarial judges. We discuss the trade-offs between efficiency and inclusivity in larger LLM evaluation ensembles and the value of the incremental LLM judge.
We release all data, code, and leaderboard at https://llm-council.com.
@@ 2 Related Work
Disagreement and dissenting voices amongst humans and LLMs. Data perspectivism [5, 30, 33] posits that differing opinions from human annotators on subjective phenomena can all be correct, reflecting their diverse cultural and demographic backgrounds, and should thus be considered true in the gold standard. In earlier language models, disagreements with human annotations were often seen as errors, hallucinations, or incompetencies. However, as LLMs exhibit more sophisticated abilities, these differences in judgment may be increasingly interpreted as valid dissenting perspectives. Notably, when humans disagreed with GPT-4, they deemed GPT-4’s judgments reasonable in 75% of cases and were willing to change their choices in 34% of cases [44]. LLMs, influenced by human data selection and algorithm design, exhibit distinct personalities [35] and biases [2, 13, 22, 29].
Using LLMs for synthetic data. LLMs can be used as a source of synthetic data for training or evaluation [38, 19, 41, 4]. By utilizing LLM(s) to generate questions, these methods can effectively mitigate the contamination concerns on static datasets and sometimes create boosts in test set variety.
LLM evaluation ensembles. Recent works have proposed using multiple agents for collaborative evaluation, mimicking a peer review process. Language-Model-as-an-Examiner [4] employs LLMs to interact with candidates through follow-up fact-oriented queries in the knowledge domain. Auto Arena [43] utilizes an LLM committee to judge competitive multi-turn interactions between LLMs. PRD [23] allows LLMs to discuss evaluations, assigning higher voting weights based on ability. PRE [11] selects a small group of reviewers to produce individual evaluations, then aggregates these evaluations through a chair. DRPE [42] uses multi-roleplayer prompting to simulate different roles with the same base model and integrates multiple outputs as votes for the final results. PoLL [36] enhances cost-effectiveness by replacing a single large judge with multiple smaller judges.
We build upon existing research by: 1) focusing on a highly subjective task where human agreement is inherently low, such as emotional intelligence, 2) using the same set of LLMs end-to-end for test set formulation, task completion, and response judging to achieve a more equitable consensus, akin to democratic social structures, and 3) engaging a large ensemble of diverse LLMs to study judging dynamics more richly. To our knowledge, this is the largest panel of LLM judges studied to date.
@@ 3 Case Study: Using the LMC to Rank LLMs on Emotional Intelligence
The Language Model Council framework consists of three stages: test set formulation, response gathering, and collective judging (Figure 1). These three stages are run sequentially and fully automated with LLM-based agents. All prompts are included in Appendix F.
@@ 3.1 Council member selection
Our selection of council member models was guided by several key considerations, including their widespread adoption within the AI community, availability of technical reports, well-supported API access, and performance on benchmarks like MMLU [20] and LMSYS [10]. We ensure a broad variety of LLMs by including models from eight different organizations across four countries, with a mix of open and closed-source models, small and large models (Figure 8).
3.2 Test set formulation and response gathering
@@
To create a compelling open-ended test set for Emotional Application (EA), we build upon the EmoBench dataset, a hand-crafted, theory-based dataset designed for this purpose [32]. EmoBench consists of 200 emotionally balanced, handcrafted questions, e.g., “Sarah found out that her younger brother is being bullied at school, but he begged her not to tell their parents.” We use the Council to transform the concise, closed-ended scenarios in EmoBench into richly described, open-ended dilemmas in the first person. See Figure 21 for the exact prompt used. Due to budget constraints, each of the 20 council members expands 5 scenarios, resulting in a dataset of 100 dilemmas, similar in scale to MT-Bench (80 questions). We manually review all expansions for EA suitability.
Relying on a single LLM to generate the entire test set, even a top performer like GPT-4o (#1 on Chatbot Arena [^1] https://chat.lmsys.org/), may introduce bias and limit perspective diversity. In a survey of 10 human respondents, 51% preferred expansions not written by GPT-4o to be included in an EA test, highlighting the subjectivity of test set construction and the need for diverse input. Inclusively constructing test sets mitigates the risk of any single LLM’s generative idiosyncrasies [1] from dominating.
Another option for inclusive test set formation is having multiple council members propose expansions for each scenario and then selecting the best through voting. This is viable but resource-intensive. Given the high quality of expansions, the incremental value of this approach is small compared to simply expanding the test set with new scenarios.
Once all 100 scenarios have been generated, each council member responds to each, resulting in 2,000 responses. In the prompt (Figure 22), we request adherence to a 250-word limit to mitigate length bias [14]. Responses exceeding this limit are truncated to the last sentence within the limit.
3.3 LLMs-as-a-jury evaluation settings
LLM rankings are based on expected win rates derived from an ELO scoring system [3, 8]. Similar to the Chatbot Arena score calculation procedure [10], we compute the Bradley-Terry (BT) coefficients [9] for better statistical estimation.
@@ 3.3 LLMs-as-a-jury evaluation settings
4-point preference scale. Based on the results of a calibration exercise (Appendix B), we query all LLMs with a temperature of 0 and use granular comparison options without ties (A>>B, A>B, B>A, B>>A) in the prompt. We use Chain-of-Thought (CoT) prompting [40] to generate answers before giving judgments.
Choice of reference respondent. Following Arena Hard [24], we use a common reference model for all pairwise battles. A dry run with 5% of the data compared GPT-4-0613 and Qwen-1.5-32B as reference models. For the main experiment, we proceed with Qwen-1.5-32B as the reference model since the battle results with Qwen-1.5-32B were more varied and resulted in significantly more separable ELO scores.
@@
Dual-sided battles. To mitigate position bias, we adopt a two-game setup, swapping model positions per query, resulting in $100*2=200$ judgments per model per judge. Following the implementation of Bradley-Terry coefficient calculations in the original codebase [^2] https://github.com/lm-sys/arena-hard-auto, inconsistent results after swapping are treated as ties and strong votes are counted as 3 separate wins.
@@
Voting aggregation functions. We consider 3 different voting functions for aggregating scores across multiple LLM judges: no aggregation (each battle judgment is equally considered); majority vote (for a given battle, we use the mode of the votes from all council members), and mean pooling (ratings are mapped to a 4-point scale (A>>B: 2, A>B: 1, B>>A: -1, B>>A: -2) and the mean of which is rounded to the nearest whole value).
@@ 3.4 Characterizing LLM judges
We leverage the many-to-many interactions between LLMs to quantify key measures of LLM judging quality in an ensemble setting.
Separability quantifies how well models can be distinguished using confidence intervals (via bootstrapping) [24]. It is measured by the percentage of model pairs with non-overlapping score confidence intervals. Higher separability indicates better model differentiation by the LLM judge. Conviction is the percentage of strong votes (e.g., A>>B or B>>A).
Consistency measures how often a judge gives consistent results when the order of two assistants is swapped. Position bias refers to an LLM favoring certain positions in pairwise comparisons [44]. Inconsistent votes are mapped to favoring either the first or second position, described in Table 10.
Agreement is measured using Cohen’s Kappa [12] between two judges’ ratings. To avoid penalizing minor order-consistent differences, granular ratings are first mapped to coarse sidewise buckets. Contrarianism is defined as the percentage of disagreement between an LLM and the Council’s majority decision.
Affinity between a judge and respondent equals the win rate the respondent model receives under the judge’s jurisdiction. Self-enhancement bias is the difference between a model’s affinity to itself and the council’s overall affinity to it. Polarization is the range of the highest and lowest assigned scores. Length bias is the $R^{2}$ of a linear regression model predicting score from average response length.

### v2

# Language Model Council: Democratically Benchmarking Foundation Models on Highly Subjective Tasks

@@ Abstract
As Large Language Models (LLMs) have become more advanced, the search for efficient and accurate evaluation methods is ongoing. Many recent evaluations use other LLMs as judges to score outputs from other LLMs, often relying on a single large model like GPT-4o. However, using a single LLM judge is prone to intra-model bias, and many tasks – such as those related to emotional intelligence, creative writing, and persuasiveness – may be too subjective for a single model to judge fairly. We introduce the Language Model Council (LMC), where a group of LLMs collaborate to create tests, respond to them, and evaluate each other’s responses in a democratic fashion. Unlike previous approaches that focus on reducing cost or bias by using a panel of smaller models, our work examines the benefits and nuances of a fully inclusive LLM evaluation system. In a detailed case study on emotional intelligence, we deploy a council of 20 recent LLMs to rank each other on open-ended responses to interpersonal conflicts. Our results show that the LMC produces rankings that are more separable and more robust, and through a user study, we show that they are more consistent with human evaluations than any individual LLM judge. Using all LLMs for judging can be costly, however, so we use Monte Carlo simulations and hand-curated sub-councils to study hypothetical council compositions and discuss the value of the incremental LLM judge.
Language Model Council: Democratically Benchmarking Foundation Models on Highly Subjective Tasks
Justin Zhao1, Flor Miriam Plaza-del-Arco2, Benjie Genchel1, Amanda Cercas Curry3 1Independent, 2Bocconi University, 3CENTAI justinxzhao@gmail.com, flor.plaza@unibocconi.it benjiegenchel@gmail.com, amanda.cercas@centai.eu https://llm-council.com
@@ 1 Introduction
Figure 1: Overview of the Language Model Council (LMC) evaluation framework. Using the same LLMs for test set formulation, task completion, and judging offers an equitable way to achieve a ranking by consensus on highly subjective tasks.
Despite recent advances in Large Language Models (LLMs), evaluating their outputs remains an open challenge. Human evaluations are challenging, time-consuming and expensive, motivating the search for automatic metrics Novikova et al. (2017); Lowe et al. (2017) and evaluation setups (e.g. Zheng et al., 2024; Li et al., 2024b; Kocmi and Federmann, 2023; Shen et al., 2023; Kim et al., 2024). Conventional model evaluations use closed-ended questions that can be checked automatically such as MMLU Hendrycks et al. (2020). These static benchmarks are at risk of contamination Ravaut et al. (2024), prompting leaderboards to keep tests private or update questions periodically (e.g. Scale AI, 2024; White et al., 2024). However, performance on these questions is often misaligned with human preferences in real-world, open-ended contexts Chiang et al. (2024). For such tasks, automatic metrics like BLEU Papineni et al. (2002), ROUGE Lin (2004), or BLEURT Sellam et al. (2020) are common. These require reference responses and fail to reflect human preferences beyond a quality threshold Freitag et al. (2020). Arena-based methods enable reference-free evaluation by judging LLMs in head-to-head comparisons. The outcome of a battle between the responses of two LLMs can be evaluated based on objective outcomes Bianchi et al. (2024), strong model judges like GPT-4 Dubois et al. (2024b); Li et al. (2024a), or human judges Li et al. (2024a), with each outcome contributing to metrics like win rates or ELO scores Chiang et al. (2024).
Strong LLM judges like GPT-4 can match both controlled and crowdsourced human preferences well, achieving over 80% agreement, the same level of agreement between humans, in a general-domain, arena-based setting Zheng et al. (2024). Unfortunately, it has also been observed that evaluator models tend to have their own biases, for example recognizing and preferring their own outputs over those of other models Dubois et al. (2024a); Panickssery et al. (2024), leading to the development of committees of LLM judges (such as Verga et al., 2024; Zhao et al., 2024).
In a landscape of diverse LLM judges, how should we aggregate each model’s vote and determine a definitive ranking among them? This paper aims to contribute a unique perspective in the pursuit of a fully decentralized approach to evaluating LLMs. When ranking LLMs on competencies that may be too subjective for a single model to judge fairly, we explore the benefits of a fully inclusive evaluation network where each model contributes equally.
Inspired by democratic organizations in human society, we introduce the Language Model Council (LMC), a framework for collective evaluation. The LMC operates in three stages: (1) collaborative test set creation by council members, (2) response generation by each LLM, and (3) evaluation of responses by the council as a collective jury.
Our main contributions are as follows:
@@
    We propose the LMC, a flexible decentralized evaluation framework that uses LLMs to rank themselves in a democratic manner, and show that our method aligns with human rankings for an open-ended emotional intelligence task.
@@
    We define and analyze key measures of LLM judging dynamics including: separability, pairwise positional consistency, agreement, and affinity, for the largest ensemble of LLM judges to date (to our knowledge).
@@
    Using all LLMs for judging can be costly, so we use Monte Carlo simulations and hand-crafted sub-councils to discuss the value of larger evaluation ensembles and the value of the incremental judge.
@@ 2 Related Work
Recently, subjective tasks and label variability in humans has received increased attention in the NLP community Alm (2011); Basile (2022); Plank (2022). LLMs also exhibit variability in their outputs as they inherit inconsistencies and biases from human data Hosking et al. (2023); Song et al. (2024); Bai et al. (2024a); Plaza-del-Arco et al. (2024); Koo et al. (2023). While LLM judges might not replicate human disagreement patterns exactly Lee et al. (2023); Dong et al. (2024), their differences in judgment are increasingly viewed as reflecting valid dissent. [^1] Notably, when humans disagreed with GPT-4, they considered its judgments reasonable in 75% of cases and sometimes revised their own answers Zheng et al. (2024).
LLM evaluation ensembles. Language-Model-as-an-Examiner Bai et al. (2024b) uses LLMs to interact with candidates through follow-up fact-oriented queries in the knowledge domain. Auto Arena Zhao et al. (2024) proposes an LLM committee to judge competitive multi-turn interactions between LLMs. PRD Li et al. (2023) allows LLMs to discuss evaluations and assigns higher voting weights based on ability. PRE Chu et al. (2024) selects a small group of reviewers to produce individual evaluations, then aggregates these evaluations through a chair model. DRPE Wu et al. (2023) uses multi-roleplayer prompting to simulate different roles with one base model, integrating them as votes for the final results. PoLL Verga et al. (2024) enhances cost-effectiveness by replacing one large judge with multiple smaller judges.
We build upon existing research by 1) focusing on a highly subjective case study on emotional intelligence where human agreement is inherently low, 2) emphasizing full inclusivity, where each LLM plays an equal role in determining the final rankings, and 3) engaging a large ensemble of diverse LLMs to study judging dynamics in greater depth. To our knowledge, this is the largest ensemble of LLM judges studied to date.
@@ 3 Case Study: Using the LMC to Rank LLMs on Emotional Intelligence
The LMC framework consists of three stages: 1) test set formulation, 2) response gathering, and 3) collective judging (Figure 1). These steps are run sequentially and are automated with LLM-based agents. [^2] We use “agents” to refer to LLMs acting simultaneously on sequential and systematic evaluation procedures. See Appendix I for all prompts.
@@ 3.1 Council member selection
Our selection of LLM council members was guided by several key considerations, including their widespread adoption within the AI community, availability of technical reports, well-supported API access, and performance on benchmarks like MMLU Hendrycks et al. (2020) and Chatbot Arena Chiang et al. (2024). We ensure a broad variety of LLMs by including models from eight different organizations and four countries, with a mix of open-source and closed-source models, small and large (Table 3 in Appendix A).
3.2 Test set formulation
@@
To create a compelling open-ended test set for Emotional Application (EA), we build upon the EmoBench dataset, a publicly available, hand-crafted, theory-based English dataset designed for EA assessment Sabour et al. (2024). EmoBench consists of 200 emotionally balanced, handcrafted scenarios, e.g., “Sarah found out that her younger brother is being bullied at school, but he begged her not to tell their parents.” We use the LMC to transform the concise, closed-ended scenarios in EmoBench into richly described, open-ended dilemmas in the first person (Figure 27). Each of the 20 council members expands five scenarios, resulting in a dataset of 100 dilemmas, similar in scale to MT-Bench (80 questions). We manually review all expansions for EA suitability. [^3] An alternative procedure is to have all council members suggest expansions for each scenario and select the best through voting. This is feasible, but given the generally high quality of expansions, the added value is limited compared to simply adding new scenarios.
Relying on a single LLM to generate the entire test set, even a top performer like GPT-4o may introduce bias and limit perspectives. In a survey of 10 human respondents, 51% of preferred expansions were not written by GPT-4o. [^4] Human respondents were asked to choose, in a series of pairwise comparisons, which expanded dilemma would be better for an emotional intelligence test. Each comparison was between a response from GPT-4 and one from a randomly chosen council member. See Appendix E for more details. Inclusively constructing test examples also mitigates the risk of any single LLM’s generative idiosyncrasies AI4Science and Quantum (2023) from dominating the test pool.
3.3 Response gathering
After generating 100 dilemmas, each council member responds to every dilemma, yielding 2,000 total responses. To standardize response lengths across council members and minimize length bias in evaluation, the prompt suggests a 250-word limit (Figure 28). Responses exceeding the limit are truncated at the nearest sentence within the limit. [^5] Sentence splits are based on standard English end punctuation (. ! ?), as the experiment is conducted in English. Despite the suggested word limit, some council members consistently generated shorter responses (Figure 1). These responses are left unchanged.
3.4 LLM-as-a-Judge evaluation settings
Arena-style pairwise comparisons with a single reference model. We adopt the pairwise comparison setup of Chatbot Arena where responses are compared in a pairwise fashion in a series of battles (Figure 30). LLM rankings are based on expected win rates derived from an ELO scoring system Bai et al. (2022); Boubdir et al. (2023). We also use Bradley-Terry (BT) coefficients (Bradley and Terry, 1952) for better statistical estimation. We use a single reference model, Qwen-1.5-32B, for all pairwise battles. For details on how the reference model was chosen, refer to Appendix D. Following Li et al. (2024a); Chiang et al. (2024), we derive confidence intervals with 100 rounds of bootstrapping.
4-point preference scale.
@@ 3.4 LLM-as-a-Judge evaluation settings
We query all LLMs with a temperature of 0 and with granular comparison options without ties (A>>B, A>B, B>A, B>>A). The reasoning behind these choices is detailed in Appendix C. We use Chain-of-Thought (CoT) prompting Wei et al. (2022) to generate discussion before giving judgments.
Exhaustive position-swapping.
@@
To minimize position bias from affecting the final ranking, we adopt a two-game setup, swapping model positions per query, resulting in $100*2=200$ judgments per model per judge. Following the implementation of BT coefficient calculation in the original codebase [^6] https://github.com/lm-sys/arena-hard-auto, inconsistent results after swapping are treated as ties and strong votes are counted as 3 separate wins.
Voting aggregations.
@@
We consider 3 different voting aggregations for consolidating scores across multiple LLM judges on a per-battle basis: majority vote (mode of all votes), mean pooling [^7] For mean pooling, we map ratings to a 4-point numeric scale (A>>B: 2, A>B: 1, B>>A: -1, B>>A: -2), take the mean rounded to the nearest whole value, and use the value corresponding to that whole number as the final rating., and no aggregation (judgments across all battles are equally considered).
@@ 3.5 Characterizing LLM judges
We leverage the many-to-many interactions between LLMs to define key judging qualities of LLMs in an ensemble setting.
Separability measures how confidently models can be distinguished in the final rankings. We adopt this metric from Li et al. (2024a), which calculates separability as the percentage of model pairs with non-overlapping confidence intervals. Higher separability means better differentiation between models. The success of score separation depends on three factors: judge discrimination, test discrimination, and participant abilities. They all contribute, and any one of them could reduce the separation to 0. In the LMC framework, potentially all three factors are governed by stochastic LLMs that are non-deterministic even with temperature=0 Chann (2023). This makes it challenging to pinpoint which component would provide the most improvement if separability is low. Our analysis of separability focuses on LLM judging, as the test set, participants, and their responses remain constant across different LMC configurations.
Pairwise Positional Consistency (PPC) measures how often a judge gives consistent results when the order of the two model responses in a pairwise comparison is swapped (Appendix C.2). For instance, if the judge ranks A>B in one comparison and B>A when the positions are swapped (a battle couplet), the judge is considered "consistent" because the preference remained the same independent of position. Position bias measures how much the LLM judge favors a specific position (either the first or second) and we define it to be $1-ppc$. On the 4-point preference scale, a battle couplet is still considered consistent as long as the relative ranking remains consistent overall — fine-grained differences such as (A>>B, B>A) or (B>>A, A>B) are tolerated. Table 10 lists all possible couplets and their consistency mappings. Conviction is the percentage of strong votes (A>>B or B>>A).
Affinity between a judge and respondent is the score the respondent model receives under the judge’s jurisdiction. Self-enhancement bias is the difference between a model’s affinity to itself and the council’s score for that model. Polarization is the range of the highest and lowest assigned scores. Length bias is the $R^{2}$ of a linear regression model predicting score from average response length.
Agreement is measured using Cohen’s Kappa Cohen (1960) between two judges’ ratings. Similar to position bias, we consider judges in agreement as long as they express the same relative preference, e.g. (A>B and A>>B) or (B>A and B>>) are in agreement. Contrarianism is measured as the disagreement between an LLM and the Council’s majority decision, reported as $1-\kappa$ [^8] Cohen’s $\kappa$ ranges from -1 to 1. Subtracting from 1 is not particularly interpretable. Applying the negative makes it so that a higher score implies more disagreement and vice versa.
3.6 Human study
To validate the LMC’s evaluations, we conduct a human study mirroring that of the LMC: human raters are asked to choose the best of two responses to a given dilemma. Note that the goal is to replicate human preferences in terms of general ranking, not to model the distribution of preferences.
Due to budget constraints, we select nine LLM council members from our pool of 20 to be rated for this study (listed in Figure 3) with variety in terms of model size, open- versus closed-source, and company origin in mind. We recruit participants via crowdsourcing on Prolific. [^9] https://www.prolific.com/ We recruited a total of 102 participants. Each response were rated by 11 participants on average, with a total of 1343 ratings. Detailed information on our recruitment process, quality control, and participant demographics is provided in Appendix E.